# GTA: Generative Trajectory Augmentation with Guidance for Offline Reinforcement Learning

### v3

@@ Abstract
Offline Reinforcement Learning (Offline RL) presents challenges of learning effective decision-making policies from static datasets without any online interactions. Data augmentation techniques, such as noise injection and data synthesizing, aim to improve Q-function approximation by smoothing the learned state-action region. However, these methods often fall short of directly improving the quality of offline datasets, leading to suboptimal results. In response, we introduce GTA, Generative Trajectory Augmentation, a novel generative data augmentation approach designed to enrich offline data by augmenting trajectories to be both high-rewarding and dynamically plausible. GTA applies a diffusion model within the data augmentation framework. GTA partially noises original trajectories and then denoises them with classifier-free guidance via conditioning on amplified return value. Our results show that GTA, as a general data augmentation strategy, enhances the performance of widely used offline RL algorithms in both dense and sparse reward settings. Furthermore, we conduct a quality analysis of data augmented by GTA and demonstrate that GTA improves the quality of the data. Our code is available at https://github.com/Jaewoopudding/GTA
@@ 1 Introduction
We introduce Generative Trajectory Augmentation (GTA), a novel approach that applies the conditional diffusion model to data augmentation, aiming to address the aforementioned limitations. GTA is designed to generate novel and high-rewarding trajectories while preserving dynamic plausibility. Our approach consists of three main stages: (1) Training a conditional diffusion model that generates trajectory-level data, (2) Augmenting the offline data via partial noising and denoising framework with amplified return guidance, and (3) Training any offline RL algorithm with augmented data.
We train a conditional diffusion model that approximates the conditional distribution of trajectory given its return. Once the diffusion model is trained, we sample trajectories from the offline dataset and partially noise the trajectory with the diffusion forward process. Then, we denoise the noised trajectory with amplified return guidance, directing trajectories to the high-rewarding region. To this end, we can orthogonally integrate data augmented by GTA into any offline RL algorithms without any modification.
Our contributions. In this paper, we introduce GTA, a novel data augmentation framework that utilizes a conditional diffusion model to generate high-rewarding, novel, and dynamically plausible data. Through extensive experiments on commonly studied benchmarks [15, 16], we demonstrate that GTA significantly improves performance across various tasks with unique challenges, such as sparse reward tasks and high-dimensional robotics tasks. We thoroughly examine the impact of core design components of GTA, trajectory generation, partial noising and denoising framework, with amplified return guidance. Furthermore, we assess the GTA augmented dataset with data quality metrics, proving its alignment with the objective of offline RL. These findings underscore the capability of GTA to efficiently augment trajectories, resulting in high-quality samples and improving the performance of offline RL algorithms.
@@ 2 Related Works
Data Augmentation for Reinforcement Learning. In Reinforcement Learning (RL), data augmentation enhances sample efficiency and improves Q-function approximation. In pixel-based RL, methods like CURL [17], RAD [11], and DrQ [18] have leveraged image augmentations such as cropping and translation to address sample-efficiency of RL. For the proprioceptive observations, S4RL [12] introduces variations into states by adding small-scale noise or adversarial gradients of Q-function under the similarity prior, that similar states yield similar rewards.
@@ 4.1 Stage A: Train Diffusion Model
Using a diffusion model to generate trajectories offers multiple advantages. First, the diffusion model leverages sequential relationships between consecutive transitions while generating trajectories, allowing it to generate dynamically plausible trajectories. Second, the diffusion model captures long-term transition dynamics, which is beneficial for environments with sparse rewards.
@@ 4.3 Stage C: Offline RL Policy Training
The final stage of GTA is utilizing high-quality trajectories generated through previous stages. The trajectories generated by GTA can seamlessly integrate with existing offline RL methods. By introducing GTA, we can orthogonally integrate the expressiveness and controllability of diffusion models with existing offline RL algorithms.
@@ 4.4 Additional Technique: Reweighted Sampling
To concentrate on samples in high-reward regions, we additionally adopt a reweighting strategy [33, 34] during sampling. This approach prioritizes the sampling of subtrajectories with higher returns. Consequently, augmented trajectories are more densely distributed in high-rewarding regions. The detailed configuration of the weight setting is provided in Section A.3.
@@ 5 Experiments
In this section, we present extensive experiments conducted across commonly studied benchmarks [15, 16] to evaluate the augmentation capability of GTA. Through experiments, we aim to address the following questions: 1) How much performance gain does GTA exhibit across various types of tasks? 2) What impacts do the design elements of GTA have on performance? 3) Are trajectories augmented via GTA high quality?
@@ 5.1 Experimental Setup
Offline RL algorithms. To demonstrate the general efficacy of GTA, we select four widely used offline RL algorithms: TD3BC [7], CQL [8], IQL [35], and MCQ [4]. For particularly challenging tasks, such as Maze2d, Antmaze, Adroit, and FrankaKitchen, we employ IQL, which provides all hyperparameters configurations across all tasks and consistently demonstrates stable performance.
@@
Algo. Aug. Cheetah-run Average
medium medium-replay medium-expert expert
None 53.3 $\pm$ 3.0 44.8 $\pm$ 3.6 50.6 $\pm$ 8.2 34.5 $\pm$ 8.3 45.8 $\pm$ 5.8
SynthER 50.4 $\pm$ 3.1 25.6 $\pm$ 2.9 36.8 $\pm$ 6.0 11.9 $\pm$ 2.4 31.2 $\pm$ 3.6
DrQ+BC GTA 53.3 $\pm$ 1.9 38.1 $\pm$ 2.5 61.0 $\pm$ 8.2 46.8 $\pm$ 8.2 49.8 $\pm$ 4.6
@@ 5.2 Benchmark Evaluations
Maze tasks. To assess GTA in sparse reward tasks, we test GTA on Maze2d and AntMaze tasks. Table 1 reveals that GTA notably enhances IQL performance on the sparse reward tasks while Synther degrades the performance. This result indicates that trajectory-level generation of GTA helps capture long-term dynamics, effectively leveraging that information while augmenting trajectories.
@@
Pixel-based Observations. We broaden the applicable area of GTA to pixel-based environments by generating data in the latent space of pretrained CNN encoder following [13]. As exhibited in Table 3, the performance boost implies the extensibility of GTA in pixel-based observations.
@@ 5.3 Ablation Studies
Figure 4: (a), (b) D4RL normalized score across different noise levels over the course of training TD3BC on halfcheetah-medium-v2 and halfcheetah-medium-expert-v2. (c) Comparison oracle reward sum of subtrajectory between conditioning strategy.
What are the benefits of trajectory-level generation? To investigate the impact of utilizing sequential relationships through trajectory-level generation, we examine performance differences as the trajectory length varied. We find that when $H=1$, the dynamic MSE becomes five times or more higher compared to trajectory-level generation. Additionally, when training TD3BC with augmented data, the normalized score increased by $21\%$ with trajectory generation compared to transition generation, where $H=1$. Detailed experiment results are in Section E.1
Is preserving original information via partial noising essential? We conduct experiments by varying the noising ratio on datasets of different optimality to explore the effect of the exploration on performance. Figure 4 illustrates that in Figure 4a halfcheetah-medium, higher noise levels enhance performance, while in Figure 4b halfcheetah-medium-expert, optimal performance occurs at a noise level of $\mu=0.25$, with declines at higher levels. This finding suggests that effective augmentation can be achieved with minimal exploration when offline data already includes high-rewarding trajectories. To this end, introducing the partial noising and denoising framework enables balancing exploration and exploitation. The impact of partial noising on preserving terminal state information is in Section F.3.
@@ 5.5 Futher Experimental Results
Larger batch size and training epochs. Since GTA offers more training data, it is worth exploring whether increasing batch size or training epochs can lead to an additional performance gain. The results in Section G.1 demonstrate that we can even enhance the performance by increasing batch size and training epochs while the baseline does not benefit from the batch size and training epochs increment.
@@
Sensitivity tests. We conduct experiments on the sensitivity of parameters $\alpha$ in Section G.2 and $\mu$ in Section G.3. The sensitivity of GTA with the amount of augmented dataset is in Section G.4. We find that using just 5% of the offline dataset with GTA outperforms the results achieved with 100% of the dataset without GTA.
@@ 6 Conclusion
We propose GTA: Generative Trajectory Augmentation, a novel generative data augmentation approach for enriching offline data with high-rewarding and dynamically plausible trajectories. GTA combines data augmentation with the diffusion sampling process by partially adding noise to original trajectories and subsequently denoising them under amplified return guidance. Our extensive experiments on 27 datasets across 8 environments exhibit considerable performance improvements in four distinct offline RL algorithms, demonstrating the versatility and effectiveness of the GTA. We show that GTA successfully creates high-quality datasets from the sub-optimal offline dataset, which leads to learning effective decision-making policy.
@@
We provide further details of our paper in the appendix. Our code implementation can be found in https://anonymous.4open.science/r/GTA/README.md
@@ E.4 Effectiveness of GTA on sequence modeling-based approaches
Unlike prior works that augment data at the transition level, GTA generates trajectories and therefore can be applied to sequence modeling approaches such as Decision Transformer [DT, 37]. For training DT, we should assign returns-to-go token for each timestep $t$, which is calculated as $\hat{R}_{t}=\sum_{k=t}^{T}r_{k}$. To model return-to-go for our generated subtrajectory, we set the return-to-go of the first timestep of the subtrajectory as $\hat{R}_{t}$, where $\hat{R}_{t}$ can be computed from the original data. For subsequent timesteps, we deduct the reward that we generated.
@@ F.4 Data Quality Table for the Figure 5
F.5 Oracle Reward Distribution of Generated Trajectories

### v4

@@ Abstract
Offline Reinforcement Learning (Offline RL) presents challenges of learning effective decision-making policies from static datasets without any online interactions. Data augmentation techniques, such as noise injection and data synthesizing, aim to improve Q-function approximation by smoothing the learned state-action region. However, these methods often fall short of directly improving the quality of offline datasets, leading to suboptimal results. In response, we introduce GTA, Generative Trajectory Augmentation, a novel generative data augmentation approach designed to enrich offline data by augmenting trajectories to be both high-rewarding and dynamically plausible. GTA applies a diffusion model within the data augmentation framework. GTA partially noises original trajectories and then denoises them with classifier-free guidance via conditioning on amplified return value. Our results show that GTA, as a general data augmentation strategy, enhances the performance of widely used offline RL algorithms across various tasks with unique challenges. Furthermore, we conduct a quality analysis of data augmented by GTA and demonstrate that GTA improves the quality of the data. Our code is available at https://github.com/Jaewoopudding/GTA
@@ 1 Introduction
We introduce Generative Trajectory Augmentation (GTA), a novel approach that applies the conditional diffusion model to data augmentation, aiming to address the aforementioned limitations. GTA is designed to generate novel and high-rewarding trajectories while minimizing degradation of dynamic plausibility. Our approach consists of three main stages: (1) Training a conditional diffusion model that generates trajectory-level data, (2) Augmenting the offline data via partial noising with diffusion forward process and denoising with amplified return guidance, and (3) Training any offline RL algorithm with augmented data.
We train a conditional diffusion model that approximates the conditional distribution of trajectory given its return. Once the diffusion model is trained, we sample trajectories from the offline dataset and partially noise the trajectory with the diffusion forward process. Then, we denoise the noised trajectory with amplified return guidance, directing trajectories to the high-rewarding region. To this end, we can orthogonally integrate data augmented by GTA into any offline RL algorithm without any modification.
Our contributions. In this paper, we introduce GTA, a novel data augmentation framework that utilizes a conditional diffusion model to generate high-rewarding, novel, and dynamically plausible data. Through extensive experiments on commonly studied benchmarks [15, 16], we demonstrate that GTA significantly improves performance across various tasks with unique challenges, such as sparse reward tasks and high-dimensional robotics tasks. We thoroughly examine the impact of core design components of GTA, trajectory generation, partial noising with diffusion forward process and denoising with amplified return guidance. Furthermore, we assess the GTA augmented dataset with data quality metrics, proving its alignment with the objective of offline RL. These findings underscore the capability of GTA to efficiently augment trajectories, resulting in high-quality samples and improving the performance of offline RL algorithms.
@@ 2 Related Works
Data Augmentation for Reinforcement Learning. In Reinforcement Learning (RL), data augmentation enhances sample efficiency and improves Q-function approximation. In pixel-based RL, methods like CURL [17], RAD [11], and DrQ [18] have leveraged image augmentations such as cropping and translation to address sample-efficiency of RL. For the proprioceptive observations, S4RL [12] introduces variations into states by adding small-scale noise or adversarial gradients of Q-function under the similarity prior, that similar states yield similar rewards. AWM [19] improves robustness of augmented data by applying simple transformations to the learned dynamics model, enabling zero-shot generalization to unseen dynamics without requiring multiple test-time rollouts.
@@ 4.1 Stage A: Train Diffusion Model
Using a diffusion model to generate trajectories offers multiple advantages. First, the diffusion model leverages sequential relationships between consecutive transitions while generating trajectories, allowing it to minimize the degradation of dynamic plausibility. Second, the diffusion model captures long-term transition dynamics, which is beneficial for environments with sparse rewards.
@@ 4.3 Stage C: Offline RL Policy Training
The final stage of GTA is utilizing high-quality trajectories generated through previous stages for policy training. The trajectories generated by GTA can seamlessly integrate with existing offline RL methods. By introducing GTA, we can orthogonally integrate the expressiveness and controllability of diffusion models with existing offline RL algorithms.
@@ 4.4 Additional Technique: Reweighted Sampling
To concentrate on samples in high-rewarding regions, we additionally adopt a reweighting strategy [34, 35] during the sampling process. This approach prioritizes the sampling of subtrajectories with higher returns. Consequently, augmented trajectories are more densely distributed in high-rewarding regions. The complete specification of our reweighting strategy, including implementation details and parameter configurations, is provided in Section A.3.
@@ 5 Experiments
In this section, we present extensive experiments conducted across commonly studied benchmarks [15, 16] to evaluate the augmentation capability of GTA. Through experiments, we aim to address the following questions: 1) How much performance gain does GTA exhibit across various types of tasks? 2) What impacts do the design elements of GTA have on performance? 3) Are trajectories augmented via GTA high quality? 4) What is the preferable $\mu$ and $\alpha$ setting for the new tasks?
@@ 5.1 Experimental Setup
Offline RL algorithms. To demonstrate the general efficacy of GTA on proprioceptive observations, we select four widely used offline RL algorithms: TD3BC [7], CQL [8], IQL [36], and MCQ [4]. For particularly challenging tasks, such as Maze2d, Antmaze, Adroit, and FrankaKitchen, we employ IQL, which provides all hyperparameter configurations across all tasks and consistently demonstrates stable performance. For pixel-based observations, we utilize DrQ+BC as the baseline algorithm, as established in the VD4RL benchmark by Lu et al. [16].
@@ 5.2 Benchmark Evaluations
Maze tasks. To assess GTA in sparse reward tasks, we test GTA on Maze2d and AntMaze tasks. Table 1 reveals that GTA notably enhances the performance of IQL policy on the sparse reward tasks while SynthER often degrades the performance. This result suggests two insights. First, the trajectory-level generation of GTA helps capture long-term dynamics, effectively leveraging that information while augmenting sparsely rewarded trajectories. Second, augmenting the dataset with high-rewarding trajectories strengthens the goal-reaching ability by enriching the demonstrations with more successful trajectories.
@@
Pixel-based Observations. We extend our approach to pixel-based observations by following the experimental setup detailed in [13]. Initially, we pretrain the policy and extract its visual encoder. Subsequently, we project offline pixel-based observations into the embedding space, followed by augmenting these embedded states with GTA. More detailed experiment setups are elaborated in Section C.4. The results presented in Table 3 show performance improvements on DrQ+BC across the environments and dataset qualities. It indicates that GTA can be further extended to pixel-based observations beyond proprioceptive observations.
@@ 5.3 Ablation Studies
Algo. Aug. Cheetah-run Walker-walk Average
medium medium-replay medium-expert expert medium medium-replay medium-expert expert
None 53.3 $\pm$ 3.0 44.8 $\pm$ 3.6 50.6 $\pm$ 8.2 34.5 $\pm$ 8.3 40.1 $\pm$3.1 13.4 $\pm$4.0 43.3 $\pm$4.1 63.3 $\pm$5.1 42.7 $\pm$4.1
SynthER 50.4 $\pm$ 3.1 25.6 $\pm$ 2.9 36.8 $\pm$ 6.0 11.9 $\pm$ 2.4 29.1 $\pm$4.4 7.6 $\pm$2.5 30.2 $\pm$2.7 28.7 $\pm$3.8 27.5 $\pm$3.3
DrQ+BC GTA 53.3 $\pm$ 1.9 38.1 $\pm$ 2.5 61.0 $\pm$ 8.2 46.8 $\pm$ 8.2 38.0 $\pm$5.0 13.6 $\pm$1.5 45.3 $\pm$5.7 67.4 $\pm$10.5 45.4 $\pm$5.7
Figure 4: (a), (b) D4RL normalized score across different noise levels over the course of training TD3BC on halfcheetah-medium-v2 and halfcheetah-medium-expert-v2. (c) Comparison oracle reward sum of subtrajectory between conditioning strategy on halfcheetah-medium-v2.
What are the benefits of trajectory-level generation? To investigate the impact of utilizing sequential relationships through trajectory-level generation, we examine performance differences as the trajectory length varied. We find that when $H=1$, the dynamic MSE becomes five times or more higher compared to trajectory-level generation with $H$ longer than 16. Additionally, when training TD3BC with augmented data, the normalized score increased by $21\%$ with trajectory generation compared to transition generation, where $H=1$. This result highlights sequential relationship between consecuive transitions in trajectory does help generation, minimizing degradation of dynamic plausibility. Detailed experiment results are in Section E.1
Is preserving original information via partial noising essential? We conduct experiments by varying the noising ratio on datasets of different optimality to explore the effect of the exploration on performance. In Figure 4a halfcheetah-medium, higher noise levels enhance performance, while in Figure 4b halfcheetah-medium-expert, optimal performance occurs at a noise level of $\mu=0.25$, with declines at higher levels. We speculate that if the original data already contains high-rewarding trajectories, even minimal modifications can be highly effective, and excessive exploration might lead to unexpected outcomes. Therefore, preserving information of original trajectory by partial noising is crucial in achieving effective data augmentation while avoiding potential negative impacts from excessive exploration. Additionally, the terminal signal, which may be very sparse, could be lost significantly if the original information is not preserved. We further conduct experiments about the impact of partial noising on preserving terminal state information in Section F.3.
@@ 5.4 Data Quality Analysis
5.5 Guideline for Selecting $\mu$ and $\alpha$
Table 4: D4RL normalized score on locomotion environments with fixed $\alpha$ and $\mu$. The experiments are conducted with TD3BC.
Algo. Halfcheetah Hopper Walker2d Average
GTA ($\mu=0.25,\alpha=1.1$) 49.22 $\pm$ 0.52 46.48 $\pm$ 0.39 94.98 $\pm$ 1.66 66.16 $\pm$ 4.89 72.86 $\pm$ 28.42 107.09 $\pm$ 2.69 85.42 $\pm$ 1.30 86.02 $\pm$ 8.98 110.67 $\pm$ 0.89 79.88 $\pm$ 3.35
Table 5: D4RL normalized score on medium and medium-replay quality locomotion environments with fixed $\alpha$ and $\mu$. The experiments are conducted with TD3BC.
Algo. Medium Medium-replay Environment Average
Halfcheetah Hopper Walker2d Halfcheetah Hopper Walker2d
None 48.42 $\pm$ 0.62 61.04 $\pm$ 3.18 84.58 $\pm$ 1.92 44.64 $\pm$ 0.71 65.69 $\pm$ 24.41 84.11 $\pm$ 4.12 64.75 $\pm$ 4.15
SynthER 49.16 $\pm$ 0.39 63.70 $\pm$ 3.69 85.43 $\pm$ 1.14 45.57 $\pm$ 0.34 78.81 $\pm$ 15.80 90.67 $\pm$ 1.56 68.89 $\pm$ 2.62
GTA ($\mu=0.5,\alpha=1.3$) 57.92 $\pm$ 0.48 68.46 $\pm$ 1.32 88.38 $\pm$ 2.70 48.23 $\pm$ 5.42 77.17 $\pm$ 22.17 91.12 $\pm$ 6.45 72.17 $\pm$ 11.23
GTA ($\mu=0.75,\alpha=1.3$) 57.85 $\pm$ 0.27 61.58 $\pm$ 5.00 87.14 $\pm$ 1.73 48.99 $\pm$ 2.16 97.26 $\pm$ 3.38 90.89 $\pm$ 3.29 79.05 $\pm$ 2.90
In this section, we propose a guideline for selecting hyperparameters based on dataset quality and environment characteristics. First, when the known information about the dataset is limited and minimizing exploration risk is essential, a low exploration and exploitation approach can enhance performance without losing stability. As illustrated in Table 4, GTA with low exploration and exploitation ($\mu=0.25,\alpha=1.1$) surpasses baseline methods in gym locomotion tasks.
However, if we know that the offline dataset is suboptimal and there exists considerable room for improvement, a higher exploration and exploitation setting can yield superior performance by uncovering a broader range of high-quality behaviors. Table 5 shows that GTA with ($\mu=0.5,\alpha=1.3$) and ($\mu=0.75,\alpha=1.3$) significantly boosts performance on suboptimal datasets. These findings indicate that while conservative parameters with low $\mu$ and $\alpha$ offer stable gains, we can elevate values of $\mu$ and $\alpha$ by leveraging prior knowledge about dataset quality to achieve superior outcomes. Notably, GTA consistently outperforms baselines across all configurations without dataset-specific hyperparameter tuning.
@@ 5.6 Futher Experimental Results
Larger batch size and training epochs. Since GTA offers more training data, it is worth exploring whether increasing batch size or training epochs can lead to additional performance gain. The results in Section G.1 demonstrate that we can enhance the performance by increasing batch size and training epochs while baselines do not benefit from the batch size and training epochs increment.
@@
Sensitivity tests. We conduct experiments on the sensitivity of parameters $\alpha$ in Section G.2 and $\mu$ in Section G.3. The sensitivity of GTA with the amount of augmented dataset is in Section G.4. Our results indicate that GTA maintains similar performance beyond one million augmented transitions, demonstrating robustness to the number of transition samples.
@@ 6 Conclusion
We propose GTA: Generative Trajectory Augmentation, a novel generative data augmentation approach for enriching offline data with high-rewarding and dynamically plausible trajectories. GTA combines data augmentation with the diffusion sampling process by partially adding noise to original trajectories and subsequently denoising them under amplified return guidance. Our extensive experiments on 31 datasets across 9 environments exhibit considerable performance improvements in four distinct offline RL algorithms, demonstrating the versatility and effectiveness of the GTA. We show that GTA successfully creates high-quality datasets from the sub-optimal offline dataset, which leads to learning effective decision-making policy.
@@
Acknowledgement
We thank the anonymous reviewers for their insightful comments and suggestions, which significantly improve our manuscript. This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2024-00410082).
@@ Acknowledgement
We provide further details of our paper in the appendix. Our code implementation can be found in https://github.com/Jaewoopudding/GTA
@@ E.4 Effectiveness of GTA on sequence modeling-based approaches
Unlike prior works that augment data at the transition level, GTA generates trajectories and therefore can be applied to sequence modeling approaches such as Decision Transformer [DT, 38]. For training DT, we should assign returns-to-go token for each timestep $t$, which is calculated as $\hat{R}_{t}=\sum_{k=t}^{T}r_{k}$. To model return-to-go for our generated subtrajectory, we set the return-to-go of the first timestep of the subtrajectory as $\hat{R}_{t}$, where $\hat{R}_{t}$ can be computed from the original data. For subsequent timesteps, we deduct the reward that we generated.
@@ F.4 Data Quality Table for the Figure 5
F.5 Comparisons of GTA with Offline Model-based RL and Diffusion Planners
In this section, we compare the performance and computational cost of GTA with offline model-based RL methods and diffusion planners. While GTA shares commonalities with both approaches as discussed in Section 2, it introduces key approaches that lead to improved performance and test-time computation efficiency.
Table 26 demonstrates superior performance of GTA compared to both model-based RL algorithms and diffusion planners in locomotion tasks. The results suggest that trajectory augmentation via diffusion models can be more effective than traditional single-transition dynamic models used in offline model-based RL. Additionally, GTA outperforms diffusion planners, which require a time-intensive sampling process during decision-making.
We also analyze the computational efficiency of GTA compared to baseline methods. As shown in Table 27, GTA exhibits significant computational advantages over diffusion planners during test-time evaluation. This advantage stems from shifting the computational burden from the sampling phase to the data preparation stage.
The model training process for both GTA and offline model-based RL algorithms [25, 27] consists of two stages: dynamic model training and policy training. GTA takes longer for dynamic model training due to its diffusion model with millions of parameters. However, offline model-based RL algorithms require significantly more time for policy training because model rollouts slow down this step. Overall, the total training time for COMBO and GTA is comparable, but GTA demonstrates superior performance, as shown in Table 26.
Table 26: Comparison with diffusion planners and offline model-based RL baselines with GTA. For baselines, we take results from their original papers. For GTA, we report score of TD3BC and CQL as base algorithms.
Dataset Type Diffusion Planners Model-based RL GTA
Diffuser DD AdaptDiffuser MOPO MOReL COMBO TD3BC CQL
halfcheetah-medium 42.8 49.1 44.2 42.3 42.1 54.2 57.84 ± 0.51 54.14 ± 0.31
halfcheetah-medium-replay 37.7 39.3 38.3 53.1 40.2 55.1 50.04 ± 0.84 51.36 ± 0.27
halfcheetah-medium-expert 88.9 90.6 89.6 63.3 53.3 90.0 93.13 ± 3.07 94.93 ± 3.71
hopper-medium 74.3 79.3 92.2 28.0 95.4 97.2 69.57 ± 4.05 74.80 ± 7.42
hopper-medium-replay 93.6 100.0 96.6 67.5 93.6 89.5 89.31 ± 16.84 98.88 ± 3.51
hopper-medium-expert 103.3 1 111.8 111.6 23.7 108.7 111.1 110.40 ± 4.04 110.90 ± 3.44
walker2d-medium 79.6 82.5 84.7 17.8 77.8 81.9 86.69 ± 0.89 80.40 ± 4.98
walker2d-medium-replay 70.6 75.0 84.4 39.0 49.8 56.0 93.82 ± 1.74 91.57 ± 5.15
walker2d-medium-expert 106.9 108.8 108.2 44.6 95.6 103.3 110.86 ± 0.34 110.44 ± 0.28
locomotion average 77.5 81.8 83.4 42.60 72.72 82.03 84.63 ± 2.20 85.27 ± 1.02
Table 27: Comparison of computational cost between Diffusion planners, offline model-based RL baselines, and GTA. The elapsed time was measured using a single RTX 3090 on the HalfCheetah-medium-v2 environment.
Diffusion Planners Model-based RL GTA
DD AdaptDiffuser MOPO COMBO TD3BC
Model Training (H) 32.0 45.3 1.7 1.7 53.0
RL Policy Training (H) - - 20.8 45.8 2.0
Synthetic Data Generation (H) - 2.4 - - 3.5
Policy Evaluation (s) 1.3 1.4 6×$\times 10^{-4}$ 7×$\times 10^{-4}$ 3×$\times 10^{-4}$
F.6 Correlation on Conditioning Value and Generated Reward
To evaluate the controllability of GTA in trajectory generation, we analyze the Pearson correlation between the sum of rewards for generated subtrajectories and the conditioning value. As shown in Figure 7, the correlation coefficients of Halfcheetah medium, Halfcheetah medium-replay and Halfcheetah medium-expert datasets are 0.55, 0.91, and 0.99, respectively. These results demonstrate that our diffusion model can effectively generate trajectories that align with the conditioned return values.
(a) halfcheetah-medium-replay
(b) halfcheetah-medium
(c) halfcheetah-medium-expert
Figure 7: Correlation on conditioning value and generated reward
F.7 Oracle Reward Distribution of Generated Trajectories